OpenStudy (anonymous):

x+y=7 x-y=-1 I don't know how to solve and graph

OpenStudy (anonymous):

Hi there. Care to work on this together?

OpenStudy (anonymous):

use algebra to isolate y x+y=7 is equal to y=7-x

OpenStudy (anonymous):

x-y=-1 is equal to -y=-1-x is equal to y=1+x

OpenStudy (anonymous):

x+y=7 x-y=-1---->y=x+1 substitute it x+x+1=7 2x=6 x=3 y=1+3=4